OpenStudy (anonymous):

I am having some issues with this problem: f(x)=x^3+2x^2-8 and g(x)=2x^2-3x+5. Evaluate the following: (f+g)(-2). I am supposed to substitute x for -2?

OpenStudy (anonymous):

no you need to add you f and g as follow:\[(f+g)(x)=f(x)+g(x)=(x^3+2x^2-8)+(2x^2-3x+5)\] \[(f+g)(x)=x^3+4x^2-3x-3\] then replace x by -2 in the expression of (f+g)(x) and simplify. You should get: \[(f+g)(-2)=-8+16+6-3=11\]